Find the different quotient for the function f(x)=4x+4 Simplify
Answer by stanbon(75887)

Find the difference quotient for the function f(x)=4x+4 Simplify
--------------------------
f(x+h) = 4(x+h)+4 = 4x+4h+4
f(x) = 4x+4
-------
Therefore f(x+h)-f(x) = 4h
---------------
And the quotient is [f(x+h)-f(x)/h = 4h/h = 4
===================================================
Cheers,
Stan H.